OpenStudy (anonymous):

Katalin drove 180 miles on her vacation. she drove an average of 1.5 times faster on the second 90 miles of her trip than she did on the first 90 miles of her trip. which expression represents the time she spent driving? let x= her speed on the first half of the trip

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

v= d/t where v = velocity, d= distance t= time solve for t so t=d/v focus on the first half of the trip, distance = 90 then focus on the other half, distance = 90 and velocity is 1.5x then add up the two t and you have your equation